# Four-Derivative Yang-Mills Couplings in Heterotic Theory through T-Duality

### v1

@@ 1 Introduction
In scenarios where the YM field is absent and field redefinitions are involved, the $O(1,1,\mathbb{Z})$ symmetry possesses the capability to determine all parameters in the minimal basis at four derivatives, which consists of 8 couplings [18], except for a single parameter. In fact, T-duality establishes 7 relations among the 8 unambiguous and essential parameters. On the other hand, if the maximal basis is employed, which encompasses 20 couplings [18], T-duality still yields 7 relations among the 20 parameters. In this case, the determination relies on a single unambiguous or essential parameter, accompanied by 12 additional arbitrary parameters. The utilization of field redefinitions enables the elimination of these arbitrary parameters [7].
In the presence of YM fields added to the NS-NS fields at the four-derivative order, the truncated T-duality establishes 24 relations among the parameters of the independent couplings. If we consider the couplings in the minimal basis, which comprises 24 couplings, the truncated T-duality fixes them to be zero. This outcome is expected since there is no string theory wherein the $H$-field possesses a YM Chern-Simons term (see eq.(2)) but lacks a Lorentz Chern-Simons term. However, when the four-derivative coupling $H\Omega$ is included, obtained through the replacement of the Lorentz Chern-Simons three-form in the $B$-field strength at the two-derivative order, the T-duality determines all 24 parameters in the minimal basis in terms of the coefficient of $H\Omega$. The NS-NS couplings derived through this approach precisely correspond to those present in the Metsaev-Tseytlin action, which were obtained using the S-matrix method [18].
@@ 4 Effective Action at the Four-Derivative Order
$+2\partial_{b}\bar{\phi}\bar{F}^{ab}+\alpha_{ij}[2(\partial_{b}\bar{\phi}-\frac{1}{4}\partial_{b}\varphi)e^{\varphi/2}V^{ab}-2(\partial_{b}\bar{\phi}+\frac{1}{4}\partial_{b}\varphi)e^{-\varphi/2}W^{ab}-e^{\varphi/2}\partial_{b}V^{ab}$
@@ 4.1 Couplings in the Minimal Scheme
${\bf S}^{(1)}$ $=$ $-\frac{\alpha^{\prime}}{4\kappa^{2}}\int d^{10}x\sqrt{-G}e^{-2\Phi}\Big{[}\frac{1}{4}F_{\alpha}{}^{\gamma kl}F^{\alpha\beta ij}F_{\beta}{}^{\delta}{}_{kl}F_{\gamma\delta ij}-\frac{1}{2}F_{\alpha}{}^{\gamma}{}_{ij}F^{\alpha\beta ij}F_{\beta}{}^{\delta kl}F_{\gamma\delta kl}$ (34)
@@
$\Delta\bar{\phi}^{(1)}$ $=$ $-\frac{1}{2}e^{\varphi}V_{ab}V^{ab}+\frac{1}{2}{e^{-\varphi}}W_{ab}W^{ab}\,,$
$\Delta G_{ab}^{(1)}$ $=$ $-2e^{\varphi}V_{a}{}^{c}V_{bc}+2{e^{-\varphi}}W_{a}{}^{c}W_{bc}\,,$
$\Delta\alpha_{ij}^{(1)}$ $=$ $-\frac{1}{2}e^{\varphi}V_{cd}V^{cd}\alpha_{ij}+\frac{1}{2}{e^{-\varphi}}W_{cd}W^{cd}\alpha_{ij}\,,$
$\Delta\bar{A}_{a}^{(1)}{}_{ij}$ $=$ $-e^{\varphi/2}\bar{H}_{ade}V^{de}\alpha_{ij}-{e^{\varphi/2}}\bar{H}_{ade}W^{de}\alpha_{ij}\,,$
$\Delta B_{ab}^{(1)}$ $=$ $4V_{[b}{}^{c}W_{a]c}+2e^{\varphi/2}\bar{F}_{[b}{}^{cij}V_{a]}{}_{c}\alpha_{ij}+2{e^{-\varphi/2}}\bar{F}_{[b}{}^{cij}W_{a]}{}_{c}\alpha_{ij}\,,$
$\Delta\varphi^{(1)}$ $=$ $-2e^{\varphi}V_{ab}V^{ab}-2{e^{-\varphi}}W_{ab}W^{ab}-2\partial_{a}\varphi\partial^{a}\varphi+2V^{ab}W_{ab}\,,$
$\Delta g_{a}^{(1)}$ $=$ $-e^{\varphi/2}\bar{H}_{abc}V^{bc}-2{e^{-\varphi/2}}\partial_{b}W_{a}{}^{b}+4{e^{-\varphi/2}}W_{ab}\partial^{b}\bar{\phi}+\frac{1}{2}e^{-\varphi/2}\bar{H}_{abc}W^{bc}-e^{\varphi/2}V_{ab}\partial^{b}\varphi$ (35)
@@
The observation that the single-trace terms of the four YM field strength are zero is consistent with the S-matrix results. However, the two-trace terms in (34) do not precisely match those obtained from S-matrix calculations [11]. This discrepancy is not an inconsistency because the two-trace terms represent ambiguous couplings that can be modified through field redefinitions. Therefore, the explicit form of the four YM field strength that is produced by the S-matrix method should appear in other schemes. To confirm the consistency of the single-trace terms with the S-matrix, we need to derive the effective action in alternative schemes. In the next subsection, we will determine the couplings in the maximal scheme.
@@ 4.2 Couplings in the Maximal Scheme
${\bf S}^{(1)}$ $=$ $-\frac{\alpha^{\prime}}{4\kappa^{2}}\int d^{10}x\sqrt{-G}e^{-2\Phi}\Big{[}\frac{1}{4}F_{\alpha}{}^{\gamma kl}F^{\alpha\beta ij}F_{\beta}{}^{\delta}{}_{kl}F_{\gamma\delta ij}+\frac{1}{2}F_{\alpha}{}^{\gamma}{}_{ij}F^{\alpha\beta ij}F_{\beta}{}^{\delta kl}F_{\gamma\delta kl}$ (39)
@@
$\Delta\bar{\phi}^{(1)}$ $=$ $-\frac{1}{2}e^{\varphi/2}\bar{F}_{abij}V^{ab}\alpha^{ij}-\frac{1}{2}{e^{-\varphi/2}}\bar{F}_{abij}W^{ab}\alpha^{ij}\,,$ (40)
$\Delta G_{ab}^{(1)}$ $=$ $-4e^{\varphi/2}\bar{F}_{\{b}{}^{cij}V_{a\}}{}_{c}\alpha_{ij}-2{e^{-\varphi/2}}\bar{F}_{\{b}{}^{cij}W_{a\}}{}_{c}\alpha_{ij}\,,$
$\Delta\alpha_{ij}^{(1)}$ $=$ $-e^{\varphi}V_{cd}V^{cd}\alpha_{ij}+{e^{-\varphi}}W_{cd}W^{cd}\alpha_{ij}+2\alpha_{ij}\partial_{c}\partial^{c}\varphi-4\alpha_{ij}\partial_{c}\varphi\partial^{c}\bar{\phi}\,,$
$\Delta\bar{A}_{a}^{(1)}{}_{ij}$ $=$ $0\,,$
$\Delta B_{ab}^{(1)}$ $=$ $4V_{[b}{}^{c}W_{a]c}+2e^{\varphi/2}\bar{F}_{[b}{}^{cij}V_{a]}{}_{c}\alpha_{ij}+2{e^{-\varphi/2}}\bar{F}_{[b}{}^{cij}W_{a]}{}_{c}\alpha_{ij}\,,$
$\Delta\varphi^{(1)}$ $=$ $-e^{\varphi}V_{ab}V^{ab}-{e^{-\varphi}}W_{ab}W^{ab}-2\partial_{a}\varphi\partial^{a}\varphi+2V^{ab}W_{ab}\,,$
$\Delta g_{a}^{(1)}$ $=$ $-e^{\varphi/2}\bar{H}_{abc}V^{bc}-2{e^{-\varphi/2}}W_{ab}\partial^{b}\varphi+\frac{1}{2}e^{-\varphi/2}\bar{H}_{abc}W^{bc}-e^{\varphi/2}V_{ab}\partial^{b}\varphi-\frac{1}{2}\bar{H}_{abc}\bar{F}^{bcij}\alpha_{ij}\,,$
@@ 5 Discussion
We have observed that the four YM field strength couplings, which are single-trace, are unambiguous because they cannot be generated through field redefinitions, Bianchi identities, or total derivative terms. However, the T-duality constraint requires their coefficients to be zero. To illustrate this more clearly, let’s consider the circular reduction of the term in equation (3.1) with coefficient $c_{2}$. By applying the reduction described in equation (4), we find that the reduced form of this term is $c_{2}\,\bar{F}_{a}{}^{c}{}_{i}{}^{k}\,\bar{F}^{abij}\,\bar{F}_{b}{}^{d}{}_{k}{}^{l}\,\bar{F}_{cdjl}+4c_{2}\,e^{\varphi/2}\,\bar{F}_{a}{}^{c}{}_{i}{}^{k}\,\bar{F}_{b}{}^{d}{}_{k}{}^{l}\,\bar{F}_{cdjl}\,V^{ab}\,\alpha^{ij}\,.$ The first term is invariant under the truncated Buscher rule stated in equation (8), while the second term transforms as $-4c_{2}\,e^{-\varphi/2}\,\bar{F}_{a}{}^{c}{}_{i}{}^{k}\,\bar{F}_{b}{}^{d}{}_{k}{}^{l}\,\bar{F}_{cdjl}\,W^{ab}\,\alpha^{ij}\,.$ No other term in equation (3.1) produces such a structure. Furthermore, since it is a single-trace term, it cannot be generated by corrections to the Buscher rules, total derivative terms, or Bianchi identities in the base space. Therefore, T-duality fixes the coefficient $c_{2}$ to be zero. A similar situation occurs for the terms in equation (3.1) with coefficients $c_{2}$, $c_{6}$, and $c_{7}$. We anticipate that a similar scenario would arise for higher-derivative couplings. Thus, we speculate that the classical effective action of the heterotic theory, at $n$ orders of derivatives, should not contain single-trace couplings of $n$ YM field strengths. This is in contrast to the classical effective action of $D$-branes, which only features single-trace couplings of YM field strengths. This distinction arises from the fact that the vertex operator in the disk-level S-matrix element appears on the boundary of the disk, and there is a Chan-Paton factor for any S-matrix element. As a result, the classical effective action of type I string theory should only contain single-trace terms. This is not inconsistent with the S-duality between the heterotic theory and type I theory [23] because the higher-derivative classical effective action of one theory at weak coupling should be transformed into the higher-derivative effective action of the other theory at strong coupling, which is not a classical effective action.
The four Yang-Mills (YM) couplings generated by T-duality in the Meissner scheme can be expressed in terms of the $t_{8}$ tensor, as shown in equation (41). On the other hand, T-duality yields the following eight-derivative coupling in the heterotic theory [24]:

### v2

@@ 1 Introduction
In scenarios where the YM field is absent and field redefinitions are involved, the $O(1,1,\mathbb{Z})$ symmetry possesses the capability to determine all parameters in the minimal basis at four derivatives, which consists of 8 couplings [18], except for a single parameter. In fact, T-duality establishes 7 relations among the 8 unambiguous and essential parameters. On the other hand, if the maximal basis is employed, which encompasses 20 couplings [18], T-duality still yields 7 relations among the 20 parameters [19]. In this case, the determination relies on a single unambiguous or essential parameter, accompanied by 12 additional arbitrary parameters. The utilization of field redefinitions enables the elimination of these arbitrary parameters.
In the presence of YM fields added to the NS-NS fields at the four-derivative order, the truncated T-duality establishes 24 relations among the parameters of the independent couplings in both minimal and maximal bases. If we consider the couplings in the minimal basis, which comprises 24 couplings, the truncated T-duality fixes them to be zero. This outcome is expected since there is no string theory wherein the $H$-field possesses a YM Chern-Simons term (see eq.(2)) but lacks a Lorentz Chern-Simons term. However, when the four-derivative coupling $H\Omega$ is included, obtained through the replacement of the Lorentz Chern-Simons three-form in the $B$-field strength at the two-derivative order, the T-duality determines all 24 parameters in the minimal basis in terms of the coefficient of $H\Omega$. The NS-NS couplings derived through this approach precisely correspond to those present in the Metsaev-Tseytlin action, which were obtained using the S-matrix method [18].
@@ 4 Effective Action at the Four-Derivative Order
$+2\partial_{b}\bar{\phi}\bar{F}^{ab}{}_{ij}+\alpha_{ij}[2(\partial_{b}\bar{\phi}-\frac{1}{4}\partial_{b}\varphi)e^{\varphi/2}V^{ab}-2(\partial_{b}\bar{\phi}+\frac{1}{4}\partial_{b}\varphi)e^{-\varphi/2}W^{ab}-e^{\varphi/2}\partial_{b}V^{ab}$
@@ 4.1 Couplings in the Minimal Scheme
${\bf S}^{(1)}$ $=$ $-\frac{2\alpha^{\prime}}{8\kappa^{2}}\int d^{10}x\sqrt{-G}e^{-2\Phi}\Big{[}\frac{1}{4}F_{\alpha}{}^{\gamma kl}F^{\alpha\beta ij}F_{\beta}{}^{\delta}{}_{kl}F_{\gamma\delta ij}-\frac{1}{2}F_{\alpha}{}^{\gamma}{}_{ij}F^{\alpha\beta ij}F_{\beta}{}^{\delta kl}F_{\gamma\delta kl}$ (34)
@@
$-8\Delta\bar{\phi}^{(1)}$ $=$ $-\frac{1}{2}e^{\varphi}V_{ab}V^{ab}+\frac{1}{2}{e^{-\varphi}}W_{ab}W^{ab}\,,$
$-8\Delta G_{ab}^{(1)}$ $=$ $-2e^{\varphi}V_{a}{}^{c}V_{bc}+2{e^{-\varphi}}W_{a}{}^{c}W_{bc}\,,$
$-8\Delta\alpha_{ij}^{(1)}$ $=$ $-\frac{1}{2}e^{\varphi}V_{cd}V^{cd}\alpha_{ij}+\frac{1}{2}{e^{-\varphi}}W_{cd}W^{cd}\alpha_{ij}\,,$
$-8\Delta\bar{A}_{a}^{(1)}{}_{ij}$ $=$ $-e^{\varphi/2}\bar{H}_{ade}V^{de}\alpha_{ij}-{e^{\varphi/2}}\bar{H}_{ade}W^{de}\alpha_{ij}\,,$
$-8\Delta B_{ab}^{(1)}$ $=$ $4V_{[b}{}^{c}W_{a]c}+2e^{\varphi/2}\bar{F}_{[b}{}^{cij}V_{a]}{}_{c}\alpha_{ij}+2{e^{-\varphi/2}}\bar{F}_{[b}{}^{cij}W_{a]}{}_{c}\alpha_{ij}\,,$
$-8\Delta\varphi^{(1)}$ $=$ $-2e^{\varphi}V_{ab}V^{ab}-2{e^{-\varphi}}W_{ab}W^{ab}-2\partial_{a}\varphi\partial^{a}\varphi+2V^{ab}W_{ab}\,,$
$-8\Delta g_{a}^{(1)}$ $=$ $-e^{\varphi/2}\bar{H}_{abc}V^{bc}-2{e^{-\varphi/2}}\partial_{b}W_{a}{}^{b}+4{e^{-\varphi/2}}W_{ab}\partial^{b}\bar{\phi}+\frac{1}{2}e^{-\varphi/2}\bar{H}_{abc}W^{bc}-e^{\varphi/2}V_{ab}\partial^{b}\varphi$ (35)
@@
The observation that the single-trace terms of the four YM field strength are zero is consistent with the S-matrix results. However, the two-trace terms in (34) do not precisely match those obtained from S-matrix calculations [11]. This discrepancy is not an inconsistency because the two-trace terms represent ambiguous couplings that can be modified through field redefinitions. Therefore, the explicit form of the four YM field strength that is produced by the S-matrix method should appear in other schemes. To confirm the consistency of the two-trace terms with the S-matrix, we need to derive the effective action in alternative schemes. In the next subsection, we will determine the couplings in the maximal scheme.
@@ 4.2 Couplings in the Maximal Scheme
${\bf S}^{(1)}$ $=$ $-\frac{2\alpha^{\prime}}{8\kappa^{2}}\int d^{10}x\sqrt{-G}e^{-2\Phi}\Big{[}\frac{1}{4}F_{\alpha}{}^{\gamma kl}F^{\alpha\beta ij}F_{\beta}{}^{\delta}{}_{kl}F_{\gamma\delta ij}+\frac{1}{2}F_{\alpha}{}^{\gamma}{}_{ij}F^{\alpha\beta ij}F_{\beta}{}^{\delta kl}F_{\gamma\delta kl}$ (39)
@@
$-8\Delta\bar{\phi}^{(1)}$ $=$ $-\frac{1}{2}e^{\varphi/2}\bar{F}_{abij}V^{ab}\alpha^{ij}-\frac{1}{2}{e^{-\varphi/2}}\bar{F}_{abij}W^{ab}\alpha^{ij}\,,$ (40)
$-8\Delta G_{ab}^{(1)}$ $=$ $-4e^{\varphi/2}\bar{F}_{\{b}{}^{cij}V_{a\}}{}_{c}\alpha_{ij}-4{e^{-\varphi/2}}\bar{F}_{\{b}{}^{cij}W_{a\}}{}_{c}\alpha_{ij}\,,$
$-8\Delta\alpha_{ij}^{(1)}$ $=$ $-e^{\varphi}V_{cd}V^{cd}\alpha_{ij}+{e^{-\varphi}}W_{cd}W^{cd}\alpha_{ij}+2\alpha_{ij}\partial_{c}\partial^{c}\varphi-4\alpha_{ij}\partial_{c}\varphi\partial^{c}\bar{\phi}\,,$
$-8\Delta\bar{A}_{a}^{(1)}{}_{ij}$ $=$ $0\,,$
$-8\Delta B_{ab}^{(1)}$ $=$ $4V_{[b}{}^{c}W_{a]c}+2e^{\varphi/2}\bar{F}_{[b}{}^{cij}V_{a]}{}_{c}\alpha_{ij}+2{e^{-\varphi/2}}\bar{F}_{[b}{}^{cij}W_{a]}{}_{c}\alpha_{ij}\,,$
$-8\Delta\varphi^{(1)}$ $=$ $-e^{\varphi}V_{ab}V^{ab}-{e^{-\varphi}}W_{ab}W^{ab}-2\partial_{a}\varphi\partial^{a}\varphi+2V^{ab}W_{ab}\,,$
$-8\Delta g_{a}^{(1)}$ $=$ $-e^{\varphi/2}\bar{H}_{abc}V^{bc}-2{e^{-\varphi/2}}W_{ab}\partial^{b}\varphi+\frac{1}{2}e^{-\varphi/2}\bar{H}_{abc}W^{bc}-e^{\varphi/2}V_{ab}\partial^{b}\varphi-\frac{1}{2}\bar{H}_{abc}\bar{F}^{bcij}\alpha_{ij}\,,$
@@ 5 Discussion
We have observed that the four YM field strength couplings, which are single-trace, are unambiguous because they cannot be generated through field redefinitions, Bianchi identities, or total derivative terms. However, the T-duality constraint requires their coefficients to be zero. To illustrate this more clearly, let’s consider the circular reduction of the term in equation (3.1) with coefficient $c_{2}$. By applying the reduction described in equation (4), we find that the reduced form of this term is $c_{2}\,\bar{F}_{a}{}^{c}{}_{i}{}^{k}\,\bar{F}^{abij}\,\bar{F}_{b}{}^{d}{}_{k}{}^{l}\,\bar{F}_{cdjl}+4c_{2}\,e^{\varphi/2}\,\bar{F}_{a}{}^{c}{}_{i}{}^{k}\,\bar{F}_{b}{}^{d}{}_{k}{}^{l}\,\bar{F}_{cdjl}\,V^{ab}\,\alpha^{ij}\,.$ The first term is invariant under the truncated Buscher rule stated in equation (8), while the second term transforms as $-4c_{2}\,e^{-\varphi/2}\,\bar{F}_{a}{}^{c}{}_{i}{}^{k}\,\bar{F}_{b}{}^{d}{}_{k}{}^{l}\,\bar{F}_{cdjl}\,W^{ab}\,\alpha^{ij}\,.$ No other term in equation (3.1) produces such a structure. Furthermore, since it is a single-trace term, it cannot be generated by corrections to the Buscher rules, total derivative terms, or Bianchi identities in the base space. Therefore, T-duality fixes the coefficient $c_{2}$ to be zero. A similar situation occurs for the terms in equation (3.1) with coefficients $c_{3}$, $c_{6}$, and $c_{7}$. We anticipate that a similar scenario would arise for higher-derivative couplings. Thus, we speculate that the classical effective action of the heterotic theory, at $n$ orders of derivatives and for $n>2$, should not contain single-trace couplings of $n$ YM field strengths. This is in contrast to the classical effective action of $D$-branes, which only features single-trace couplings of YM field strengths. This distinction arises from the fact that the vertex operator in the disk-level S-matrix element appears on the boundary of the disk, and there is a Chan-Paton factor for any S-matrix element. As a result, the classical effective action of type I string theory should only contain single-trace terms. This is not inconsistent with the S-duality between the heterotic theory and type I theory [24] because the higher-derivative classical effective action of one theory at weak coupling should be transformed into the higher-derivative effective action of the other theory at strong coupling, which is not a classical effective action.
The four YM couplings generated by T-duality in the Meissner scheme can be expressed in terms of the $t_{8}$ tensor, as shown in equation (41). On the other hand, T-duality yields the following eight-derivative coupling in the heterotic theory [25]: